OpenStudy (anonymous):

The formula for the amount with simple interest is A = P(1 + rt). Find t if A = 168, P = 150, and r = 0.08. A. 1.5 B. 2 C. 2.4 D. 3

OpenStudy (anonymous):

A = P(1+rt) => 1+rt = A/P => rt = A/P - 1 => t = (A/P - 1)/r

OpenStudy (anonymous):

Now, it is a simply matter of plugging in the values.

OpenStudy (angela210793):